OpenStudy (anonymous):

tell which rate is greater 1 1/6 miles in 2 min and 4 sec or 1 and 3/16 mi. in 1 min and 55 sec

OpenStudy (anonymous):

1 1/6=7/6miles

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

here, rate=(7/6)/(2*60+4)=0.0094 miles per second

OpenStudy (anonymous):

another rate = (3/16)/(1*60+55)=0.0016

OpenStudy (anonymous):

Which one do u think is the answer

OpenStudy (anonymous):

i meant which is greater lol like which is bigger

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

1 1/6 miles in 2 min and 4 sec is greater